# A Survey on Occupancy Perception for Autonomous Driving: The Information Fusion Perspective

### v2

@@ 4.2.2 Inference Speed
Table 5: Inference speed analysis of 3D occupancy perception on the Occ3D-nuScenes [35] dataset. $\dagger$ indicates data from FullySparse [107]. $\ddagger$ means data from FastOcc [82]. R-50 represents ResNet50 [38]. TRT denotes acceleration using the TensorRT SDK [143].
@@
Recent studies on 3D occupancy perception [82, 107] have begun to consider not only perception accuracy but also its inference speed. According to the data provided by FastOcc [82] and FullySparse [107], we sort out the inference speeds of 3D occupancy methods, and also report their running platforms, input image sizes, backbone architectures, and occupancy accuracy on the Occ3D-nuScenes dataset, as depicted in Tab. 5.
@@ 5.1 Occupancy-based Applications in Autonomous Driving
3D occupancy perception enables a comprehensive understanding of the 3D world and supports various tasks in autonomous driving. Existing occupancy-based applications include segmentation, detection, dynamic perception, and world models. (1) Segmentation: Semantic occupancy perception can essentially be regarded as a 3D semantic segmentation task. (2) Detection: OccupancyM3D [5] and SOGDet [144] are two occupancy-based works that implement 3D object detection. OccupancyM3D first learns occupancy to enhance 3D features, which are then used for 3D detection. SOGDet develops two concurrent tasks: semantic occupancy prediction and 3D object detection, training these tasks simultaneously for mutual enhancement. (3) Dynamic perception: Cam4DOcc [10] predicts foreground flow in 3D space from the perspective of occupancy and achieves an understanding of 3D environmental changes. (4) World models: OccNet [8] quantifies physical 3D scenes into semantic occupancy and trains a shared occupancy descriptor. This descriptor is fed to various task heads to implement driving tasks. For example, the motion planning head outputs planning trajectories for ego vehicles, and the detection head outputs 3D boxes of objects. Similarly, UniScene [60], Occupancy-MAE [99] and DriveWorld [7] pre-train occupancy to extract a universal representation of the 3D scene, and then fine-tune various downstream tasks, such as 3D object detection, online mapping, multi-object tracking, motion forecasting, occupancy prediction, and planning.
However, existing occupancy-based applications primarily focus on the perception level, and less on the decision-making level. Given that 3D occupancy is more consistent with the 3D physical world than other perception manners (e.g., bird’s-eye view perception and perspective-view perception), we believe that 3D occupancy holds opportunities for broader applications in autonomous driving. At the perception level, it could improve the accuracy of existing place recognition [145], pedestrian detection [146, 147], accident prediction [148], and lane line segmentation [149]. At the decision-making level, it could help safer driving decisions [150] and navigation [151, 152], and provide 3D explainability for driving behaviors.
@@ 5.2 Deployment Efficiency
Currently, some real-time efforts on occupancy tasks have been attempted. For instance, Hou et al. [82] proposed a solution, FastOcc, to accelerate the prediction inference speed based on the adjustment of the input resolution, view transformation module, and prediction head. Liu et al. [107] proposed SparseOcc, a sparse occupancy network without any dense 3-D features, to minimize the computational costs based on the sparse convolution layers and mask-guided sparse sampling. Tang et al. [90] proposed to adopt the sparse latent representation instead of TPV representation and sparse interpolation operation to avoid information loss and reduce computational complexity. However, the above-mentioned approaches are still some way away from real-time deployment for the adoption of autonomous driving systems.
@@ 5.3 Robust 3D Occupancy Perception
However, research into robust 3D occupancy is limited, primarily due to the scarcity of datasets. Recently, the ICRA 2024 RoboDrive Challenge [154] provides imperfect scenarios for studying robust 3D occupancy perception. We consider that related works in robust BEV perception [155, 156, 157, 158, 46, 47] could inspire the study of robust occupancy perception. M-BEV [156] proposes random masking and reconstructing camera views to enhance robustness under various missing camera cases. GKT [157] employs coarse projection to achieve a robust BEV representation. In most scenarios involving natural damage, multi-modal models [158, 46, 47] outperform single-modal models by the complementary nature of multi-modal input. Besides, in 3D LiDAR perception, Robo3D [159] distills knowledge from a teacher model with complete point clouds to a student model with imperfect input, thereby enhancing the student model’s robustness. Based on these works, approaching to robust 3D occupancy perception could include, but is not limited to, robust data representation, multiple modalities, network architecture, and learning strategies.
@@ 5.4 Generalized 3D Occupancy Perception
3D labels are costly, and large-scale 3D annotations for the real world are impractical. The generalization capabilities of existing networks trained on limited 3D-labelled datasets have not been extensively studied. To get rid of dependence on 3D labels, self-supervised learning represents a potential pathway toward generalized 3D occupancy perception. It learns occupancy perception from a broad range of unlabelled images. However, the performance of current self-supervised occupancy perception [87, 37, 91, 30] is poor. On the Occ3D-nuScene dataset (see Tab. 4), the top accuracy of self-supervised methods is inferior to that of strongly-supervised methods by a large margin. Moreover, current self-supervised methods require training and evaluation with more data. Thus, enhancing self-supervised generalized 3D occupancy perception is an important future research direction.
@@ Acknowledgment
The research work was conducted in the JC STEM Lab of Machine Learning and Computer Vision funded by The Hong Kong Jockey Club Charities Trust.

### v3

@@ 4.2.2 Inference Speed
Table 6: Inference speed analysis of 3D occupancy perception on the Occ3D-nuScenes [36] dataset. $\dagger$ indicates data from SparseOcc [118]. $\ddagger$ means data from FastOcc [82]. R-50 represents ResNet50 [39]. TRT denotes acceleration using the TensorRT SDK [159].
@@
Recent studies on 3D occupancy perception [82, 118] have begun to consider not only perception accuracy but also its inference speed. According to the data provided by FastOcc [82] and SparseOcc [118], we sort out the inference speeds of 3D occupancy methods, and also report their running platforms, input image sizes, backbone architectures, and occupancy accuracy on the Occ3D-nuScenes dataset, as depicted in Tab. 6.
@@ 5 Challenges and Opportunities
In this section, we explore the challenges and opportunities of 3D occupancy perception in autonomous driving. Occupancy, as a geometric and semantic representation of the 3D world, can facilitate various autonomous driving tasks. We discuss both existing and prospective applications of 3D occupancy, demonstrating its potential in the field of autonomous driving. Furthermore, we discuss the deployment efficiency of occupancy perception on edge devices, the necessity for robustness in complex real-world driving environments, and the path toward generalized 3D occupancy perception.
@@ 5.1 Occupancy-based Applications in Autonomous Driving
3D occupancy perception enables a comprehensive understanding of the 3D world and supports various tasks in autonomous driving. Existing occupancy-based applications include segmentation, detection, dynamic perception, world models, and autonomous driving algorithm frameworks. (1) Segmentation: Semantic occupancy perception can essentially be regarded as a 3D semantic segmentation task. (2) Detection: OccupancyM3D [5] and SOGDet [6] are two occupancy-based works that implement 3D object detection. OccupancyM3D first learns occupancy to enhance 3D features, which are then used for 3D detection. SOGDet develops two concurrent tasks: semantic occupancy prediction and 3D object detection, training these tasks simultaneously for mutual enhancement. (3) Dynamic perception: Its goal is to capture dynamic objects and their motion in the surrounding environment, in the form of predicting occupancy flows for dynamic objects. Strongly-supervised Cam4DOcc [10] and self-supervised LOF [160] have demonstrated potential in occupancy flow prediction. (4) World model: It simulates and forecasts the future state of the surrounding environment by observing current and historical data [161]. Pioneering works, according to input observation data, can be divided into semantic occupancy sequence-based world models (e.g., OccWorld [162] and OccSora [163]), point cloud sequence-based world models (e.g., SCSF [108], UnO [164], PCF [165]), and multi-camera image sequence-based world models (e.g., DriveWorld [7] and Cam4DOcc [10]). However, these works still perform poorly in high-quality long-term forecasting. (5) Autonomous driving algorithm framework: It integrates different sensor inputs into a unified occupancy representation, then applies the occupancy representation to a wide span of driving tasks, such as 3D object detection, online mapping, multi-object tracking, motion prediction, occupancy prediction, and motion planning. Related works include OccNet [8], DriveWorld [7], and UniScene [61].
However, existing occupancy-based applications primarily focus on the perception level, and less on the decision-making level. Given that 3D occupancy is more consistent with the 3D physical world than other perception manners (e.g., bird’s-eye view perception and perspective-view perception), we believe that 3D occupancy holds opportunities for broader applications in autonomous driving. At the perception level, it could improve the accuracy of existing place recognition [166, 167], pedestrian detection [168, 169], accident prediction [170], and lane line segmentation [171]. At the decision-making level, it could help safer driving decisions [172] and navigation [173, 174], and provide 3D explainability for driving behaviors.
@@ 5.2 Deployment Efficiency
Currently, some real-time and deployment-friendly efforts on occupancy tasks have been attempted. For instance, Hou et al. [82] proposed a solution, FastOcc, to accelerate prediction inference speed by adjusting the input resolution, view transformation module, and prediction head. Zhang et al. [175] further lightweighted FlashOcc by decomposing its occupancy network and binarizing it with binarized convolutions. Liu et al. [118] proposed SparseOcc, a sparse occupancy network without any dense 3-D features, to minimize computational costs using sparse convolution layers and mask-guided sparse sampling. Tang et al. [90] proposed to adopt sparse latent representations and sparse interpolation operations to avoid information loss and reduce computational complexity. Additionally, Huang et al. recently proposed GaussianFormer [154], which utilizes a series of 3D Gaussians to represent sparse interest regions in space. GaussianFormer optimizes the geometric and semantic properties of the 3D Gaussians, corresponding to the semantic occupancy of the interest regions. It achieves comparable accuracy to state-of-the-art methods using only 17.8%-24.8% of their memory consumption. However, the above-mentioned approaches are still some way from practical deployment in autonomous driving systems. A deployment-efficient occupancy method requires superiority in real-time processing, lightweight design, and accuracy simultaneously.
@@ 5.3 Robust 3D Occupancy Perception
However, research on robust 3D occupancy is limited, primarily due to the scarcity of datasets. Recently, the ICRA 2024 RoboDrive Challenge [177] provides imperfect scenarios for studying robust 3D occupancy perception.
In terms of network architecture and scene representation, we consider that related works on robust BEV perception [47, 48, 178, 179, 180, 181] could inspire developing robust occupancy perception. M-BEV [179] proposes a masked view reconstruction module to enhance robustness under various missing camera cases. GKT [180] employs coarse projection to achieve robust BEV representation. In terms of sensor modality, radar can penetrate small particles such as raindrops, fog, and snowflakes in adverse weather conditions, thus providing reliable detection capability. Radar-centric RadarOcc [182] achieves robust occupancy perception with imaging radar, which not only inherits the robustness of mmWave radar in all lighting and weather conditions, but also has higher vertical resolution than mmWave radar. RadarOcc has demonstrated more accurate 3D occupancy prediction than LiDAR-centric and vision-centric methods in adverse weather. Besides, in most damage scenarios involving natural factors, multi-modal models [47, 48, 181] usually outperform single-modal models, benefiting from the complementary nature of multi-modal inputs. In terms of training strategies, Robo3D [97] distills knowledge from a teacher model with complete point clouds to a student model with imperfect input, enhancing the student model’s robustness. Therefore, based on these works, approaches to robust 3D occupancy perception could include, but are not limited to, robust scene representation, multiple modalities, network design, and learning strategies.
@@ 5.4 Generalized 3D Occupancy Perception
Although more accurate 3D labels mean higher occupancy prediction performance [183], 3D labels are costly and large-scale 3D annotations for the real world are impractical. The generalization capabilities of existing networks trained on limited 3D-labeled datasets have not been extensively studied. To get rid of dependence on 3D labels, self-supervised learning represents a potential pathway toward generalized 3D occupancy perception. It learns occupancy perception from a broad range of unlabelled images. However, the performance of current self-supervised occupancy perception [31, 38, 87, 91] is poor. On the Occ3D-nuScene dataset (see Tab. 4), the top accuracy of self-supervised methods is inferior to that of strongly-supervised methods by a large margin. Moreover, current self-supervised methods require training and evaluation with more data. Thus, enhancing self-supervised generalized 3D occupancy perception is an important future research direction.
@@ Acknowledgment
The research work was conducted in the JC STEM Lab of Machine Learning and Computer Vision funded by The Hong Kong Jockey Club Charities Trust and was partially supported by the Research Grants Council of the Hong Kong SAR, China (Project No. PolyU 15215824).